# An accurate vector optically pumped magnetometer with microwave-driven Rabi frequency measurements

### v1

@@ 5 MPE Calibration and Long-Term Drift
An essential step to performing vector magnetometry with the Rabi measurements is to calibrate the microwave field parameters. Because of the microwave dependence of the cavity modes, 12 phasors $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ must be calibrated, corresponding to 60 microwave parameters. We fit each $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ using Eq. (3) over $N=12$ DC field directions by minimizing the following cost function
@@
Figure 3: (a) The sequence of $\vec{B}_{\text{dc}}$ directions $(\alpha,\beta)$ consisting of random directions (black circles) interspersed with repeated directions $(0\degree,57\degree)$ and $(135\degree,57\degree)$. While all $\vec{B}_{\text{dc}}$ directions in this sequence are utilized in the MPE calibrations, vector measurements are only evaluated at the random $(\alpha,\beta)$. Conversely, the repeated field directions (gray and white circles) are used to consistently monitor microwave drifts. To avoid overfitting, the microwave phasor $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ at each random $\vec{B}_{\text{dc}}$ direction in this sequence is determined from the weighted average of the two phasors $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{prev}}$ and $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{sub}}$ extracted from the MPE calibrations that consist of the $N=12$ directions immediately before and after the random $(\alpha,\beta)$ direction. The calibration length $N$ is limited by the finite time $\Delta t=2.5$ s that is dominated by measurement dead time reserved for cavity heating. (b) Measured MPEs. The nT-scale variation of the microwave field components arising from the frequency dependence of the cavity modes are not discernible in this plot. (c) Residuals of the $\sigma^{+}$ generalized Rabi frequency $(\tilde{\Omega}_{m=1}^{m^{\prime}=2})_{p=2}$ at random DC magnetic field directions with and without MPE recalibration. (d) Drift of the $\mathcal{B}_{x}$ microwave component evaluated at different $\nu_{\mu\text{w}}=\overline{\nu}_{m}^{m^{\prime}}$. Here, MPE 2 measurements are used as a representative example in (c,d).
@@
Over 37.5 minutes, we observe nT-scale drifts of the microwave field components and mrad-scale drifts in the phases as exemplified for MPE 2 in Fig. 3(d). The microwave field drift is dominated by the contraction of the microwave cavity modes due to a few °$\degree$°C cooling of the microwave cavity. Details on monitoring the temperature drift from the Rabi dephasing rate and accounting for temperature-dependent shifts in $\nu_{\text{bg}}$ are discussed in Supplementary 1. For these calibrations, we found that it was sufficient to recalibrate only the relative phase $\delta\phi=\phi_{y}-\phi_{x}$ by leaving $\phi_{x}$ fixed after an initial calibration instead of resolving for both $\phi_{x}$ and $\phi_{y}$. We note that most of this 37.5 minute duration is spent heating the cavity and not taking measurements. One path to eliminate this in future work is to employ continuous laser heating [62] that does not perturb the atomic measurements.
To assess the calibration accuracy, we show residuals of $\sigma^{+}$ generalized Rabi frequencies $(\tilde{\Omega}_{m=1}^{m^{\prime}=2})_{p=2}$ against the $\delta\lambda_{m=1}^{m^{\prime}=2}$ model utilizing recalibrated MPE 2 parameters in Fig. 3(c) as a representative example. The residuals are only evaluated over random magnetic field orientations. Over 37.5 minutes these residuals exhibit a standard deviation of $\sigma=13$ Hz, which corresponds to $\sigma=0.8\text{ nT}$ fluctuations for the spherical microwave component $\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)}$. These fluctuations are approximately three times greater than the anticipated $\sigma=5$ Hz ($0.3$ nT), a figure expected due to the limited calibration length of $N=12$ (see Supplement 1). This discrepancy suggests that the systematic errors in the calibrated MPE parameters are estimated to be $\delta\mathcal{B}_{x}=0.4$ nT, $\delta\mathcal{B}_{y}=0.3$ nT, $\delta\mathcal{B}_{z}=0.2$ nT, and $\delta(\phi_{y}-\phi_{x})=0.3$ mrad.
@@ 6 Simultaneous Precession and Rabi (SPaR)
Figure 4: Rabi rate detection using SPaR. (a) Directional dependence of Rabi oscillation and SPaR signal amplitude. This plot shows the fast Fourier transform (FFT) of SPaR time-domain signals like the one shown in Fig. 2(a). (b) Energy-level diagram in the rotating frame near the $\sigma^{+}$ transition frequency. Coupling between microwave dressed states (green) with Zeeman sublevels (black) are indicated by double-sided arrows that indicate $\nu_{\text{hi}}$ (long, dashed) and $\nu_{\text{low}}$ (short, solid) frequency components. (c) Generalized Rabi frequency measurements using Rabi oscillations (green) and SPaR (black). SPaR measurements are shifted by the values in parenthesis. (d) Measured SPaR spectrum for the $\sigma^{+}$ transition. Vertical lines indicate $\nu_{\text{low}}$ and $\nu_{\text{hi}}$ precessional frequencies calculated from the eigenstates of $H$.
@@ 7 Vector Magnetometry Evaluation
As a sensitivity comparison, the vector component sensitivities reported in Table 1 are at the same 1 nT$/\sqrt{\text{Hz}}$ level reached in the Helium SWARM mission  [31]. While this is not a direct comparison due to a variety of apparatus differences such as the smaller cell volume employed in this work, it demonstrates comparable vector sensitivity to high accuracy OPM implementations. In future work, the implementation of complete 3D control over the microwave field, rather than using planar MPEs, would permit the strategic selection of MPEs to achieve optimal vector sensitivity at all field directions.

### v2

@@ 5 MPE Calibration and Long-Term Drift
An essential step to performing vector magnetometry with the Rabi measurements is to calibrate the microwave field parameters. Because of the microwave frequency dependence of the cavity modes, 12 phasors $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ must be calibrated, corresponding to 60 microwave parameters. We fit each $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ using Eq. (3) over $N=12$ DC field directions by minimizing the following cost function
@@
Figure 3: (a) The sequence of $\vec{B}_{\text{dc}}$ directions $(\alpha,\beta)$ consisting of random directions (black circles) interspersed with repeated directions $(0\degree,57\degree)$ and $(135\degree,57\degree)$. While all $\vec{B}_{\text{dc}}$ directions in this sequence are utilized in the MPE calibrations, vector measurements are only evaluated at the random $(\alpha,\beta)$. Conversely, the repeated field directions (gray and white circles) are used to consistently monitor microwave drifts. To avoid overfitting, the microwave phasor $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p}$ at each random $\vec{B}_{\text{dc}}$ direction in this sequence is determined from the weighted average of the two phasors $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{prev}}$ and $(\vec{\mathcal{B}}_{m}^{m^{\prime}})_{p,\text{sub}}$ extracted from the MPE calibrations that consist of the $N=12$ directions immediately before and after the random $(\alpha,\beta)$ direction. The calibration length $N$ is limited by the finite time $\Delta t=2.5$ s that is dominated by measurement dead time reserved for cavity heating. (b) Measured MPEs (solid lines) and their projections onto the coordinate planes (dotted lines). The nT-scale variation of the microwave field components arising from the frequency dependence of the cavity modes are not discernible in this plot. (c) Residuals of the $\sigma^{+}$ generalized Rabi frequency $(\tilde{\Omega}_{m=1}^{m^{\prime}=2})_{p=2}$ at random DC magnetic field directions with and without MPE recalibration. (d) Drift of the $\mathcal{B}_{x}$ microwave component evaluated at different $\nu_{\mu\text{w}}=\overline{\nu}_{m}^{m^{\prime}}$. Here, MPE 2 measurements are used as a representative example in (c,d).
@@
Over 37.5 minutes, we observe nT-scale drifts of the microwave field components and mrad-scale drifts in the phases as exemplified for MPE 2 in Fig. 3(d). The microwave field drift is dominated by the contraction of the microwave cavity modes due to a few °$\degree$°C cooling of the microwave cavity. Details on monitoring the temperature drift from the Rabi dephasing rate and accounting for temperature-dependent shifts in $\nu_{\text{bg}}$ are discussed in Supplementary 1. For these calibrations, we found that it was sufficient to recalibrate only the relative phase $\delta\phi=\phi_{y}-\phi_{x}$ by leaving $\phi_{x}$ fixed after an initial calibration instead of resolving for both $\phi_{x}$ and $\phi_{y}$. We note that most of this 37.5 minute duration is spent heating the cavity and not taking measurements. Calibrating all MPE parameters across 12 random $\vec{B}_{\text{dc}}$ directions currently requires 132 seconds. A promising approach to eliminate this duration in future work involves the use of continuous laser heating [62], which would not interfere with atomic measurements. By disregarding technical deadtime, the calibration time could be reduced to 1.8 seconds, as the total measurement duration for Rabi oscillation data is 150 ms. This duration could be further shortened to a few hundred milliseconds by reducing the MPE model parameters, for instance, by eliminating the frequency dependence of the cavity modes with improved microwave engineering. Shortening the calibration time would also help minimize systematic errors caused by potential drifts in the DC background magnetic field during MPE calibration.
To assess the calibration accuracy, we show residuals of $\sigma^{+}$ generalized Rabi frequencies $(\tilde{\Omega}_{m=1}^{m^{\prime}=2})_{p=2}$ against the $\delta\lambda_{m=1}^{m^{\prime}=2}$ model utilizing recalibrated MPE 2 parameters in Fig. 3(c) as a representative example. The residuals are only evaluated over random magnetic field orientations. Over 37.5 minutes these residuals exhibit a standard deviation of $\sigma=13$ Hz, which corresponds to $\sigma=0.8\text{ nT}$ fluctuations for the spherical microwave component $\mathcal{B}_{\sigma^{+}}^{(\alpha,\beta)}$. These fluctuations are approximately three times greater than the anticipated $\sigma=5$ Hz ($0.3$ nT), a figure expected due to the limited calibration length of $N=12$ (see Supplement 1). This discrepancy suggests that the systematic errors in the calibrated MPE parameters are estimated to be $\delta\mathcal{B}_{x}=0.4$ nT, $\delta\mathcal{B}_{y}=0.3$ nT, $\delta\mathcal{B}_{z}=0.2$ nT, and $\delta(\phi_{y}-\phi_{x})=0.3$ mrad. These calibration errors are consistent with the directional accuracies deduced in the vector magnetometry evaluation in Sec. 7.
@@ 6 Simultaneous Precession and Rabi (SPaR)
Figure 4: Rabi rate detection using SPaR. (a) Directional dependence of Rabi oscillation and SPaR signal amplitude. This plot shows the fast Fourier transform (FFT) of SPaR time-domain signals like the one shown in Fig. 2(a). (b) Energy-level diagram in the rotating frame near the $\sigma^{+}$ transition frequency. Coupling between microwave dressed states (green) with Zeeman sublevels (black) are indicated by double-sided arrows that indicate $\nu_{\text{hi}}$ (long, dashed) and $\nu_{\text{low}}$ (short, solid) frequency components. (c) Generalized Rabi frequency measurements using Rabi oscillations (green) and SPaR (black). SPaR measurements are shifted by the values in parenthesis. (d) Measured SPaR spectrum for the $\sigma^{+}$ transition. The top plot shows the real part of the SPaR FFT at the $\sigma^{+}$ transition frequency $\nu_{\mu\text{w}}=\overline{\nu}_{m=1}^{m^{\prime}=2}$, both with and without the broadened Larmor resonances occuring at $\nu_{\text{L}}$. A green dashed line represents a fit applied to the SPaR FFT after removing the Larmor resonances. The bottom plot shows the absolute value of the SPaR FFT over a range of microwave frequencies centered about $\nu_{\mu\text{w}}=\overline{\nu}_{m=1}^{m^{\prime}=2}$, with the broadened Larmor resonances retained. Vertical lines indicate $\nu_{\text{low}}$, $\nu_{\text{hi}}$, and $\nu_{\text{L}}$ precessional frequencies calculated from the eigenstates of $H$.
@@ 7 Vector Magnetometry Evaluation
As a sensitivity and bandwidth comparison, the vector component sensitivities reported in Table 1 are at the same 1 nT$/\sqrt{\text{Hz}}$ level reached in the Helium SWARM mission  [31]. With a measurement duration of $t_{m}=150$ ms, the vector magnetometer would achieve a bandwidth of 3.3 Hz if not for the currently implemented deadtime required for cavity heating and eddy current dissipation. For perspective, the vector mode of the ESA Swarm 4He magnetometer, operating in near-Earth orbit, has a bandwidth of 0.4 Hz [27]. While this is not a direct comparison due to a variety of apparatus differences such as the smaller cell volume employed in this work, it demonstrates comparable vector sensitivity and bandwidth to high accuracy OPM implementations. Future advancements could include implementing full 3D control of the microwave field, moving beyond planar MPEs. This would enable the strategic selection of MPEs to optimize vector sensitivity across all field directions. Additionally, as outlined in Sec. 5, continuous laser heating [62] offers a promising approach to eliminating deadtime associated with cavity heating. Replacing the microwave cavity with a dielectric resonator platform is another viable path to mitigate eddy current effects while reducing the sensor’s overall size and weight.